User:Anthony Salvagno/Notebook/Research/2010/05/11/Return of the Tweezers with Power Spectrum
I made a double sample of beads. Each sample chamber had a different bead type: one had the .97um beads and the other had the .51um beads. I diluted both sets of beads 1:400 and sonicated just prior to use. I flowed 10ul BGB as a precursor and then flowed 10ul of each dilution of beads. I sealed the chambers with nail polish and let sit for 10 min.
For the tweezer, I cleaned the objective and the condenser and put the sample in. I aligned the condenser. We then tried to do some power spectrum analysis.
Power Spectrum Notes:
- For the software:
- I want to make sure QPD Math and Auto Guess are off
- Even though the fan is not touching the microscope, the vibrations are still disrupting the detector. We need to shut it off. I need to find out where the new live feed software (because it has a camera shut off button) is so I can put a shortcut on the PC desktop.
About the microscopy:
- Oil is fucking annoying. When looking at samples, at first everything was blurry. Some time later, everything was clear but then again even later it was blurry again.
- The detector is more stable now. Pranav and Andy rigged something to hold the detector still. Works like a charm.
- Pranav also fixed the detector movement issue. We were having problems moving the detector around. The micrometers that were on it would run out of room and we wouldn't be able to center the beam on the QPD. This is now not a problem.
- We managed to trap both big and small beads
- Trapping only happened at higher power (around 200mW). Last time I did this, I could trap at 45mW just fine.
- We messed with the first lens in the telescope and this may have messed stuff up. I'm sick of changing one little thing on the tweezers and everything being royally fucked up afterward. It's really annoying.
- The reason for the alteration was that I wanted to position the focus of the trap so that it was near the surface of the glass.